OpenStudy (anonymous):

if you write 14 to the power of f in expanded form how many factors will there be

OpenStudy (anonymous):

i mean if you write 14 to the 6 power in expanded form how many factors are there

zepdrix (zepdrix):

Oh :)

zepdrix (zepdrix):

\[\Large\bf\sf 3^2\quad=\quad 3\cdot 3\]\[\Large\bf\sf 4^3\quad=\quad 4\cdot4\cdot4\]The power tells us how many factors we should have.

OpenStudy (anonymous):

what is the sum of 2 to the power of 4 and seven to the power of 2 write in standard form

zepdrix (zepdrix):

\[\Large\bf\sf 2^4+7^2\]Order of Operations tells us to do our `exponents` before we add. \[\Large\bf\sf 2^4\quad=\quad 2\cdot2\cdot2\cdot2\quad=\quad ?\]\[\Large\bf\sf 7^2\quad=\quad 7\cdot7\quad=\quad ?\]

OpenStudy (anonymous):

i get it thank you